OpenStudy (anonymous):

need steps to solve for x 3/2 (x-1) - (x+2)/3= 1 - (x+11)/6

OpenStudy (anonymous):

(3x-3)/2 - (x+2)/3 = 1 - (x+11)/6 multiply by 6: 3(3x-3) -2(x+2) = 6 - (x+11) 9x - 9 -2x -4 = 6 -x -11 8x = 8 x = 1

OpenStudy (anonymous):

uhm how did you get 3(3x-3) and -2(x+2) did you split the fraction or something ?